How many kJ of heat are needed to raise the temperature of 50mg of water by 10.64 degrees C.

Given:

Concentration NaOH = 0.05moles

Molar Enthalpy for dissolution of NaOH = -44517.8J

Molar Mass of NaOH = 40g/mol

Volume of water = 1.71 x10^15L
Answer #1

Mass of water = 50 mg = 50×10-3 g

Heat needed to raise temperature of 50 mg water by 10.64 degree c

= Mass of water (g) × specific heat of water × temperature change

= 50 × 10-3 ( g) × 4.184 (J/goc) × 10.64 (0c)

= 44.5 J

= 4.45 ×10-2 KJ.